# SimuHome: A Temporal- and Environment-Aware Benchmark for Smart Home LLM Agents

### v1

@@ Abstract
Large Language Model (LLM) agents excel at multi-step, tool-augmented tasks. However, smart homes introduce distinct challenges, requiring agents to handle latent user intents, temporal dependencies, device constraints, scheduling, and more. The main bottlenecks for developing smart home agents with such capabilities include the lack of a realistic simulation environment where agents can interact with devices and observe the results, as well as a challenging benchmark to evaluate them. To address this, we introduce SimuHome, a time-accelerated home environment that simulates smart devices, supports API calls, and reflects changes in environmental variables. By building the simulator on the Matter protocol [^1] https://csa-iot.org/all-solutions/matter/—the global industry standard for smart home communication—SimuHome provides a high-fidelity environment, and agents validated in SimuHome can be deployed on real Matter-compliant devices with minimal adaptation. We provide a challenging benchmark of 600 episodes across twelve user query types that require the aforementioned capabilities. Our evaluation of 11 agents under a unified ReAct framework reveals that while models perform well on simple tasks, they struggle with latent intent inference, state verification, and especially temporal scheduling. Even the top-performing model, GPT-4.1, reaches only 54% success rate. These findings highlight a critical need for methods that can reliably verify the current state via tools before acting and coordinate time-dependent actions. [^2] We will release our code and dataset upon publication of the paper.
@@ 1 Introduction
Our first contribution is a smart home simulator, SimuHome (Figure 1). SimuHome is a time-accelerated smart home environment that accommodates various room layouts, environmental variables (e.g., temperature, illuminance), and smart devices. Agents can call APIs to operate devices (e.g., set the AC to 25 degrees). Devices are simulated with internal constraints checked (e.g., the AC must be turned on to set its temperature), and the results affect the environment (e.g., the room temperature gradually drops to 25 degrees over 10 minutes). Notably, SimuHome implements Matter, a broadly adopted smart-home interoperability standard. As a result, the attributes and constraints within devices are high fidelity. Moreover, agents trained and verified in SimuHome can run on real Matter-compliant devices with minimal adaptation. SimuHome also enables controlled experiments in a cheap and fast way. It allows unlimited experimentation, including stress-testing rare edge cases and counterfactual scenarios, while strict reproducibility ensures fair comparisons and iterative validation across models. Although beyond the scope of our work, it can also support model training through reinforcement learning.
@@
We evaluate 11 LLM agents under a unified ReAct (Yao et al., 2023) setup across 600 episodes with feasible and infeasible variants, scoring feasible tasks by simulator state comparisons and assessing infeasible logic checks with validated LLM judges. Models handle simple retrieval and explicit device control comparatively well. They often fail to infer latent intent or to verify current device and environment states before acting. Temporal scheduling is the most challenging area: contradiction blindness and mishandling are common, and even GPT-4.1 reaches only 54% success rate. These results motivate methods that verify the current state via tools before acting and that plan and coordinate time-dependent actions reliably in dynamic home environments.
@@ 3.2 Simulator Architecture and Operation
Smart Home Environment. A home is a configurable environment composed of one or more rooms, each containing a custom set of devices and four environmental variables: temperature, illuminance, humidity, and air quality. To enable realistic scenarios, the environment includes both devices that directly influence environmental variables (e.g., an air conditioner) and those with multi-stage operational cycles (e.g., a washing machine). In total, we model 17 distinct device types. A full list of these devices can be found in Appendix A.4.
Real-Time State Update Mechanism. The core of the simulation is the Aggregator module, which models the dynamic impact of device operations on the environment. At each tick, the Aggregator calculates the combined influence of all active devices on their relevant environmental factors. For example, temperature is affected by air conditioners and heat pumps, illuminance by lights, humidity by humidifiers/dehumidifiers, and air quality by air purifiers. The magnitude of this influence is cumulative; it scales with the number of active devices and their specific settings (e.g., the fan speed of an air conditioner). This mechanism ensures that the environment responds realistically to agent actions. The detailed update equations for the Aggregator are provided in Appendix A.12.
Agent-Simulator Interface. The agent interacts with the simulator by invoking a set of 13 tools. The structure of these tools mirrors Matter’s modular approach to defining device capabilities. Detailed tool specifications are provided in Appendix A.2.
@@ 4.1 Query Types
We define twelve query types that commonly arise in user queries within smart home environments. These are designed to evaluate an agent’s abilities in device control, environmental variable queries such as temperature and illuminance, implicit intent inference, and temporal coordination with three sub-types. Each type is paired with an infeasible scenario to test the agent’s capacity for logical consistency and constraint handling, yielding a total of 12 categories. See Appendix A.1 for examples of infeasible scenarios corresponding to each query type.
@@ 4.2 Episode Generation
STEP2: Goals and Required Actions. A goal defines the desired final state of specific devices or environmental variables that the agent must achieve. The generation process, which varies by query type (see Appendix A.5), is designed to ensure all goals are logically consistent. For instance, as illustrated for QT3 in Figure 2 (Step 3), a device goal is created by sampling from a pre-defined set of valid states (e.g., onoff: on, fanspeed: 40%). Each of these state sets is constructed to inherently satisfy the device’s internal dependencies. Required Actions are a sequence of tool calls that an agent must perform. This ensures the agent’s subsequent actions are based on up-to-date information gathered from the environment. For example, before attempting to change an air purifier’s fan speed, the agent is required to first invoke the tool get_room_devices(utility_room) to confirm the device’s existence. An episode is marked as successful only if the agent both satisfies the goal and its tool call history contains all required actions.
STEP3: Query Synthesis. In general, a user’s natural language query embodies a goal to be achieved, and the clarity of this goal is essential for an accurate evaluation of the agent’s success. Therefore, we first defined a verifiable goal for the agent to accomplish and subsequently synthesized a natural language query based on it. We then used GPT-5-mini (OpenAI, 2025b) to synthesize the natural language queries from these predefined goals. To ensure each query accurately reflected its predefined goal, two graduate students researching tool agents independently reviewed the entire dataset. Their inter-annotator agreement, measured using Cohen’s $\kappa$ coefficient (Cohen, 1960), was 0.92 for identifying queries that required correction. This demonstrates that the validation procedure for our dataset is highly consistent and reliable, suggesting that the benchmark data is composed of high-quality natural language queries.
@@ 5.1 Evaluation Methods
Simulator-based Evaluation. Simulator-based evaluation is essential for episodes that target physical state changes because outcomes must be assessed objectively and reliably. At the end of each episode, the simulator automatically verifies the final states of all relevant devices and environmental variables, and compares them with the goal defined for that episode. In the QT4-Feasible episode shown in Figure 3, after the agent completes all its actions, the simulator accelerates time to 19:00, when the laundry cycle finishes. Subsequently, the goal state—the washer being off and the kitchen light being on—is compared with the simulator’s final state to determine success. This direct state comparison yields a clear and fully automated success criterion and enables fair model-to-model comparisons under the same conditions.
LLM-judge-based Evaluation. We employ an LLM-based judge for episodes where success depends not on physical state changes, but on the agent’s final natural-language response. This provides an objective and effective complement to simulator-based evaluation. In our procedure, the judge is provided with the episode goal and the user query, along with a concise description of any infeasible conditions that must be verified. We also provide the agent’s full ReAct trajectory so that the judge can verify whether the final answer is supported by a coherent and sufficient chain of reasoning. For example, in the QT4-Infeasible episode shown in Figure 3, the LLM-judge assesses the agent’s explanation of a scheduling conflict. The judge scores only the final user-facing answer while using the trajectory to confirm consistency. For each case, we asked the judge three times and adopted the consistent outcome to ensure reliability (Taubenfeld et al., 2025). Detailed prompt templates and rubrics are available in Appendix A.9. To validate our LLM-judges, we confirmed that their assessments achieved substantial agreement (Cohen’s $\kappa$ = 0.826) with human evaluations, which themselves showed very high inter-rater reliability (see Appendix A.6 for the full analysis).
5.2 Experiments
Experimental Setup. We quantitatively compare models’ inference, planning, and tool-calling performance across the 12 query types defined in §4.1. The evaluation covers 11 models spanning open-source (Llama, Qwen, Gemma) and closed-source (Gemini, GPT) families, with reproducibility details provided in Appendix A.7. We employ the ReAct framework (Yao et al., 2023) for all experiments, enabling the models to generate both reasoning traces and actions in a step-by-step manner within our smart home simulation environment. The specific agent prompt used for the ReAct framework is provided in Appendix A.8.
5.3 Main Results
@@
Model FJ IFJ FS IFJ FS IFJ FS IFJ FS IFJ FS IFJ
Llama-4-Scout $58$ $42$ 22 $22$ $24$ $34$ 44 44 22 22 22 0
Llama-4-Maverick 96 $78$ $52$ $36$ $88$ $74$ $22$ $14$ $18$ $10$ 32 88
Qwen3-32B $82$ $66$ 62 $30$ $52$ $68$ $18$ $14$ $14$ 88 $16$ 66
@@
Gemma-3-12B-it $78$ $38$ $14$ $32$ $32$ $24$ 22 0 0 0 0 0
Gemma-3-27B-it $80$ $48$ $54$ $24$ $48$ $44$ 44 22 $10$ 88 0 66
Gemini-2.5-Flash-Lite $78$ $60$ $44$ 50 $50$ $50$ 88 $34$ $10$ $16$ $16$ 2020
Gemini-2.5-Flash $92$ $86$ $66$ $54$ $82$ $74$ $22$ $44$ 40 32 $12$ 32
GPT-4.1-nano $58$ $42$ 66 $12$ $30$ $16$ 22 66 66 0 0 0
GPT-4.1-mini 96 $76$ 62 $28$ $64$ 76 26 40 40 2020 $10$ $28$
@@
Table 1 reports performance across all query types (QT1-QT4) under both feasible (F) and infeasible (IF) conditions. Several consistent patterns emerge.
Environment perception (QT1). QT1-F episodes are largely solved: Most frontier models exceeded 85%. In QT1-IF, we observe a success rate degradation across models, reflecting a consistent challenge in detecting requests based on false premises, such as references to non-existent devices.
Explicit vs. implicit device control (QT2 vs. QT3). Across both feasible and infeasible settings, most models handle explicit commands (QT3) substantially better than implicit requests (QT2). In feasible episodes, leading models exceed 84% on QT3, whereas QT2 plateaus at 62-66%. In feasible settings, models consistently perform better when users issue explicit commands (QT3) compared to implicit requests (QT2). In infeasible settings, however, the contrast becomes more pronounced: models can often detect nonexistent devices in QT3-IF with relatively high success rate, but they struggle far more when the device exists yet the goal is unattainable due to operational constraints. For example, an air conditioner already operating at maximum fan speed cannot lower the temperature further. This indicates that reasoning over such constraint-driven infeasibility poses a greater challenge for current models.
Temporal reasoning (QT4). QT4 queries, which require agents to handle scheduling, coordination, and temporal consistency, were the most challenging across both feasible and infeasible settings. GPT-4.1 achieved 50% on QT4-1, 46% on QT4-2, and 34% on QT4-3, outperforming peers but still far from robust. In infeasible cases with contradictory or impossible timing, performance degraded sharply: even GPT-4.1 and Gemini-2.5-Flash peaked at 44% (QT4-1-IF) and 34% (QT4-2-IF), with Qwen3-235B-A22B delivering notable performance at 48% on QT4-3-IF. These results highlight that contradiction detection and temporal validation remain open challenges.
Model-level comparison. Closed-source models (GPT-4.1, Gemini-2.5-Flash) consistently led overall, but none achieved stable performance across both feasible and infeasible tasks. Mid-sized open-source models such as Qwen3-235B-A22B and Gemini-2.5-Flash-Lite produced notable results on certain query types—for instance, Qwen3-235B-A22B reached 48% on QT4-3-IF—but their performance was uneven across query types. Smaller models (Gemma-3-12B-it, GPT-4.1-nano) remained limited overall, particularly on temporally demanding tasks.
Table 2: Error taxonomy. Detailed descriptions and examples are provided in Appendix A.10.
@@
Figure 4: Error type distributions of GPT-4.1 on QT2 and QT4. (a) and (b) show the distributions for feasible episodes, while (c) and (d) present the distributions for infeasible episodes.
6 Analysis
6.1 Error Analysis
We define eight error types to analyze agent failures: five for feasible episodes (EP, II, DC, AP, TR) and three for infeasible episodes (CM, CB, LJ). Error taxonomy is provided in Table 2. Our analysis centers on GPT-4.1, the best-performing model.
Figure 4 summarizes the error type distributions for GPT-4.1 across feasible and infeasible episodes. For feasible episodes, figure (a) and (b) show error distribution in QT2 and QT4. In QT2 (indirect requests), failures were dominated by Device Control (DC, 71%), where the model issued heuristic guesses instead of using the correct API. Intent Inference (II) errors (11%) also appeared, reflecting difficulty in mapping vague complaints such as “The room is too hot” to the appropriate device action. QT4 (temporal scheduling) exhibited a more diverse mix: DC (40%), Temporal Reasoning (TR, 25%), and Action Planning (AP, 19%) all contributed substantially, alongside smaller II errors (11%). These distributions show that multi-step temporal reasoning requires coordinating multiple skills simultaneously, making it substantially harder than direct execution tasks.
@@ 6.1 Error Analysis
6.2 Role of Tool Feedback

### v2

@@ Abstract
Large Language Model (LLM) agents excel at multi-step, tool-augmented tasks. However, smart homes introduce distinct challenges, requiring agents to handle latent user intents, temporal dependencies, device constraints, scheduling, and more. The main bottlenecks for developing smart home agents with such capabilities include the lack of a realistic simulation environment where agents can interact with devices and observe the results, as well as a challenging benchmark to evaluate them. To address this, we introduce SimuHome, a time-accelerated home environment that simulates smart devices, supports API calls, and reflects changes in environmental variables. By building the simulator on the Matter protocol [^1] https://csa-iot.org/all-solutions/matter/, the global industry standard for smart home communication, SimuHome provides a high-fidelity environment, and agents validated in SimuHome can be deployed on real Matter-compliant devices with minimal adaptation. We provide a challenging benchmark of 600 episodes across twelve user query types that require the aforementioned capabilities. Our evaluation of 16 agents under a unified ReAct framework reveals distinct capabilities and limitations across models. Models under 7B parameters exhibited negligible performance across all query types. Even GPT-4.1, the best-performing standard model, struggled with implicit intent inference, state verification, and particularly temporal scheduling. While reasoning models such as GPT-5.1 consistently outperformed standard models on every query type, they required over three times the average inference time, which can be prohibitive for real-time smart home applications. This highlights a critical trade-off between task performance and real-world practicality. [^2] We will release our code and dataset upon publication of the paper.
@@ 1 Introduction
Our first contribution is a smart home simulator, SimuHome (Figure 1). SimuHome is a time-accelerated smart home environment that accommodates various room layouts, environmental variables (e.g., temperature, illuminance), and smart devices. Agents can call APIs to operate devices (e.g., set the AC to 25 degrees). Devices are simulated with internal constraints checked (e.g., the AC must be turned on to set its temperature), and the results affect the environment (e.g., the room temperature gradually drops to 25 degrees over 10 minutes). Notably, SimuHome implements Matter, a broadly adopted smart-home interoperability standard. As a result, the attributes and constraints within devices are high fidelity. Moreover, agents trained and verified in SimuHome can run on real Matter-compliant devices with minimal adaptation.
SimuHome also enables controlled experiments in a cheap and fast way. It allows unlimited experimentation, including stress-testing rare edge cases and counterfactual scenarios, while strict reproducibility ensures fair comparisons and iterative validation across models. Although beyond the scope of our work, it can also support model training through reinforcement learning.
@@
We evaluate 16 LLM agents under a unified ReAct (Yao et al., 2023) setup across 600 episodes with feasible and infeasible variants, scoring feasible tasks by simulator state comparisons and assessing infeasible tasks with validated LLM judges. Standard models handle simple retrieval and explicit device control well, but struggle to infer latent intent and verify current states before acting. Among all query types, temporal scheduling proves most challenging. Even GPT-4.1, the best-performing standard model, achieves only 12–50% success. Reasoning models such as GPT-5.1 improve substantially (44–100%), but their threefold inference overhead limits their use in real-time applications. This motivates developing efficient methods that verify system state via tools and reliably coordinate time-dependent actions.
@@ 3.2 Simulator Architecture and Operation
Smart Home Environment. A home is a configurable environment composed of one or more rooms, each containing a custom set of devices and four environmental variables: temperature, illuminance, humidity, and air quality. To enable realistic scenarios, the environment includes both devices that directly influence environmental variables (e.g., an air conditioner) and those with multi-stage operational cycles (e.g., a washing machine). In total, we model 17 distinct device types. A full list of these devices can be found in Appendix LABEL:app:device_list.
Real-Time State Update Mechanism. The core of the simulation is the Aggregator module, which models the dynamic impact of device operations on the environment. At each tick, the Aggregator calculates the combined influence of all active devices on their relevant environmental factors. For example, temperature is affected by air conditioners and heat pumps, illuminance by lights, humidity by humidifiers/dehumidifiers, and air quality by air purifiers. The magnitude of this influence is cumulative; it scales with the number of active devices and their specific settings (e.g., the fan speed of an air conditioner). This mechanism ensures that the environment responds realistically to agent actions. The detailed update equations for the Aggregator are provided in Appendix LABEL:app:aggregator_equation.
Agent-Simulator Interface. The agent interacts with the simulator by invoking a set of 17 tools. The structure of these tools mirrors Matter’s modular approach to defining device capabilities. Detailed tool specifications are provided in Appendix B.
@@ 4.1 Query Types
We define twelve query types that commonly arise in user queries within smart home environments. These are designed to evaluate an agent’s abilities in device control, environmental variable queries such as temperature and illuminance, implicit intent inference, and temporal coordination with three sub-types. Each type is paired with an infeasible scenario to test the agent’s capacity for logical consistency and constraint handling, yielding a total of 12 categories. See Appendix A for examples of infeasible scenarios corresponding to each query type.
@@ 4.2 Episode Generation
STEP2: Goals and Required Actions. A goal defines the desired final state of specific devices or environmental variables that the agent must achieve. The generation process, which varies by query type (see Appendix LABEL:app:goal_examples), is designed to ensure all goals are logically consistent. For instance, as illustrated for QT3 in Figure 2 (Step 3), a device goal is created by sampling from a pre-defined set of valid states (e.g., onoff: on, fanspeed: 40%). Each of these state sets is constructed to inherently satisfy the device’s internal dependencies. Required Actions are a sequence of tool calls that an agent must perform. This ensures the agent’s subsequent actions are based on up-to-date information gathered from the environment. For example, before attempting to change an air purifier’s fan speed, the agent is required to first invoke the tool get_room_devices(utility_room) to confirm the device’s existence. An episode is marked as successful only if the agent both satisfies the goal and its tool call history contains all required actions.
STEP3: Query Synthesis. In general, a user’s natural language query embodies a goal to be achieved, and the clarity of this goal is essential for an accurate evaluation of the agent’s success. Therefore, we first defined a verifiable goal for the agent to accomplish and subsequently synthesized a natural language query based on it. We then used GPT-5-mini (OpenAI, 2025c) to synthesize the natural language queries from these predefined goals. To ensure each query accurately reflected its predefined goal, two graduate students researching tool agents independently reviewed the entire dataset. Their inter-annotator agreement, measured using Cohen’s $\kappa$ coefficient (Cohen, 1960), was 0.92 for identifying queries that required correction. This demonstrates that the validation procedure for our dataset is highly consistent and reliable, suggesting that the benchmark data is composed of high-quality natural language queries.
@@ 4.3 Evaluation Methods
Simulator-based Evaluation. Simulator-based evaluation is essential for episodes that target physical state changes because outcomes must be assessed objectively and reliably. At the end of each episode, the simulator automatically verifies the final states of all relevant devices and environmental variables, and compares them with the goal defined for that episode. In the QT4-Feasible episode shown in Figure 3, after the agent completes all its actions, the simulator accelerates time to 19:00, when the laundry cycle finishes. The goal state (washer off, kitchen light on) is then compared with the simulator’s final state to determine success. This direct state comparison enables fully automated and fair model-to-model comparisons. We apply simulator-based evaluation to all feasible episodes in QT2, QT3, and QT4, which involve physical state changes in the home environment.
LLM-judge-based Evaluation. We employ an LLM-based judge for episodes where success depends on the agent’s final natural-language response rather than physical state changes. The judge receives the episode goal, user query, and the agent’s full reasoning trajectory, along with a description of any infeasible conditions that must be verified. This allows the judge to assess whether the final answer is supported by coherent reasoning. For example, in the QT4-Infeasible episode shown in Figure 3, the LLM-judge evaluates the agent’s explanation of a scheduling conflict.
We apply LLM-judge-based evaluation to all infeasible episodes (QT1-IF through QT4-IF), which require assessing whether the agent correctly identifies and explains constraint violations. Additionally, QT1-Feasible also uses LLM-judge evaluation because success depends on providing accurate information in natural language rather than changing device states. For reliability, we query the judge three times per case and adopt the consistent outcome (Taubenfeld et al., 2025). Our LLM-judges achieved substantial agreement (Cohen’s $\kappa$ = 0.826) with human evaluations (see Appendix LABEL:app:llm_judge_validation). Detailed prompt templates are in Appendix LABEL:app:llm_judge_prompt.
5 Experiments
@@
Models FJ IFJ FS IFJ FS IFJ FS IFJ FS IFJ FS IFJ
Open Source Large Language Models (<7B)
Llama3.2-1B-it 0 0 0 0 0 0 0 0 0 0 0 0
Llama3.2-3B-it 10 12 0 2 4 0 2 0 2 0 0 0
Gemma3-4B-it 44 32 12 10 28 8 0 0 2 0 0 4
Open Source Large Language Models (Standard)
Llama4-Scout 58 42 2 22 24 34 4 4 2 2 2 0
Llama4-Maverick 96 78 52 36 88 74 22 14 18 10 32 8
Qwen3-32B 82 66 62 30 52 68 18 14 14 8 16 6
@@
Gemma3-12B-it 78 38 14 32 32 24 2 0 0 0 0 0
Gemma3-27B-it 80 48 54 24 48 44 4 2 10 8 0 6
Closed Source Large Language Models (Standard)
Gemini2.5-Flash-Lite 78 60 44 50 50 50 8 34 10 16 16 20
Gemini2.5-Flash 92 86 66 54 82 74 22 44 40 32 12 32
GPT-4.1-nano 58 42 6 12 30 16 2 6 6 0 0 0
GPT-4.1-mini 96 76 62 28 64 76 26 40 40 20 10 28
@@
Closed Source Large Language Models (with Reasoning)
Gemini2.5-Pro 96 78 60 56 76 72 44 94 60 76 46 50
GPT-5.1 100 94 80 50 86 92 60 100 72 92 56 44
Experimental Setup. We evaluate 16 models across the 12 query types defined in §4.1, spanning four categories: open-source models under 7B parameters, open-source standard models, closed-source standard models, and closed-source reasoning models. All experiments use the ReAct framework (Yao et al., 2023), enabling step-by-step reasoning and action generation. Reproducibility details and agent prompts are in Appendix LABEL:app:experimental_setup and LABEL:app:react_prompt.
5.1 Main Results
Table 1 presents the performance across all query types (QT1-QT4). We analyze the results by categorizing models into three groups based on scale and reasoning capability.
Lightweight Models. Models under 7B exhibit severe limitations across all query types. For instance, Llama3.2-1B-it achieves 0% across all tasks, while Gemma3-4B-it shows marginal success on basic tasks such as information retrieval (QT1-F: 44%) and explicit commands (QT3-F: 28%). However, these models completely fail on complex reasoning tasks, with near-zero success rates on implicit intent inference (QT2) and temporal reasoning (QT4), indicating insufficient reasoning capabilities for our benchmark.
Standard Models. This group comprises closed-source models (e.g., GPT-4.1, Gemini2.5-Flash) and large open-source models (e.g., Llama4-Maverick). Even GPT-4.1, the best-performing standard model, struggles with implicit intent inference (QT2) and temporal reasoning (QT4). GPT-4.1 achieves only 44% on QT2-F, compared to 84% on explicit device control (QT3-F). This gap highlights the difficulty of interpreting ambiguous user queries. Performance on temporal reasoning (QT4) plateaus at approximately 30–50% for feasible episodes and degrades further in infeasible scenarios (12–34%), revealing limited ability to detect temporal contradictions.
Reasoning Models. Advanced reasoning models (Gemini2.5-Pro, GPT-5.1) demonstrate substantial improvements over standard models, particularly in temporal reasoning and implicit intent inference. On temporal reasoning tasks (QT4-F), GPT-5.1 achieves 56–72% compared to GPT-4.1’s 34–50%. The improvement is especially pronounced in infeasible episodes, where GPT-5.1 reaches 100% on QT4-1-IF and 92% on QT4-2-IF, compared to GPT-4.1’s 12% and 34%. This demonstrates stronger capability in detecting temporal contradictions. Beyond temporal reasoning, GPT-5.1 also achieves 80% on QT2-F compared to GPT-4.1’s 44%, indicating improved interpretation of latent user intents.
5.2 Analysis
Table 2: Error taxonomy. Detailed descriptions and examples are provided in Appendix LABEL:app:error_taxonomy_details.
@@ 5.2 Analysis
Figure 4: Error type distributions of GPT-4.1 on QT2 and QT4.
5.2.1 Error Analysis.
We define eight error types to analyze agent failures: five for feasible episodes (EP, II, DC, AP, TR) and three for infeasible episodes (CM, CB, LJ). Error taxonomy is provided in Table 2. Our analysis centers on GPT-4.1, the best-performing model. Figure 4 summarizes the error type distributions for GPT-4.1 across feasible and infeasible episodes. For feasible episodes, figure (a) and (b) show error distribution in QT2 and QT4.
In QT2 (indirect requests), failures were dominated by Device Control (DC, 71%), where the model issued heuristic guesses instead of using the correct API. Intent Inference (II) errors (11%) also appeared, reflecting difficulty in mapping vague complaints such as “The room is too hot” to the appropriate device action.
QT4 (temporal scheduling) exhibited a more diverse mix: DC (40%), Temporal Reasoning (TR, 25%), and Action Planning (AP, 19%) all contributed substantially, alongside smaller II errors (11%). These distributions show that multi-step temporal reasoning requires coordinating multiple skills simultaneously, making it substantially harder than direct execution tasks.
@@
Table 3: Average episode completion time (seconds) across query types. While reasoning models achieve high accuracy, they incur significant latency costs compared to the standard model (GPT-4.1).
QT1 QT2 QT3 QT4-1 QT4-2 QT4-3
Model F IF F IF F IF F IF F IF F IF
Gemini-2.5-Pro 24.1 22.4 57.5 48.8 66.1 27.8 74.0 12.5 57.7 37.0 53.7 53.1
GPT-5.1 35.7 38.4 109.4 99.6 78.6 54.3 121.1 13.5 135.1 76.0 112.7 111.0
GPT-4.1 8.3 7.8 23.6 20.2 22.9 9.4 26.6 12.3 28.7 23.7 29.7 25.9